OpenStudy (anonymous):

How do you do this equation? 8(3y-8)=

OpenStudy (anonymous):

whats to the right of =?

OpenStudy (anonymous):

Nothing

OpenStudy (anonymous):

that's how the equation is

OpenStudy (anonymous):

How do you do it?

jhonyy9 (jhonyy9):

24y-64= so if equal 0 --- 24y-64=0 --- 24y=64 --- / 8 --- 3y=8 --- y=8/3

OpenStudy (anonymous):

Ok so explain this step by step jhonny9

jhonyy9 (jhonyy9):

multiply 8 with 3y what will be 24y and 8 with -8 equal -64 hence you get 24y-64 for you can resolve it you need equal with zero - write like one equation and after you get 24y=64 where you see that it's possible simplify with 8 and you get this result hope that you understand now

OpenStudy (anonymous):

So the answer is Y=8/3?s Did you move the 3 underneath the 8?

jhonyy9 (jhonyy9):

after you have simplified the both sides with 8 you get 3y=8 - so now for you get y - you need to divide the both sides with 3 and hence you get y=8/3